Now the friends try a homework problem. Two forces of the same magnitude 3.00 x 10 N but with opposite direction are appliedFumiko now challenges Sierra to work a homework problem. A long steel bar of length L; = 2.000 m is clamped at one end and is​​​​​​​
Answer #1

1.

Shear Modulus, T = ?? Shear Stress Shear Strain shear eshrar

Oshear ==   Eshear

\therefore \tau = \frac{\frac{F}{A}}{\frac{\Delta x}{h}} = \frac{Fh}{A\Delta x}   Fh 4.1 =

Given,

Shear Modulus,  T = 5 * 1010 N/m Force Applied F = 3 * 107N

Area, A= 0,020m2 h = 0.10m

Fh 3* 10 * 0.1 Ar = À = 0.02 * 5 * 1010 o = 0.003m = 3mm ===> Answer

2.

Young Modulus, Y = - Linear Stress Linear Strain

\sigma = \frac{F}{A}e = \frac{\Delta L}{L}

Y =\Rightarrow A = \frac{FL}{Y\Delta L}

Given,

Youngs Modulus, Y = 20 * 100 N/m Applied Force, F = 2 * 106N

Elongation , 62000 = 7V Length, L = 2m

FL 2* 100 * 2 YAL - 20 * 1010 000 = 25.31 * 10-4 m2